January 25, 2018 Drift Chamber meeting
- Instructions for Bluejeans meeting connection
- Meeting ID: 290664653
- Status and current run analysis:
Participants: Naomi (CMU) Eugene, Beni, Alex B., Simon, Sergey, Alex A., and Lubomir (JLab).
Status and current run analysis
- We discussed the settings for higher current running (up to 300nA) tomorrow (Fri 01/26/18). Eugene: we need the same conditions as for the normal running and if the chambers trip we will lower the beam current. Naomi proposed to lower the CDC HV of the inner HV sector (4 inner rings) from 2125 to 2025 so that they have current similar to the outer rings. Eugene wants to postpone this studies for later time when we have enough statistics and have dedicated studies of the DAQ rates and detector performance at higher fluxes.
- Naomi and Beni looked at the fADC thresholds - no need to be changed.
- Naomi has issues with her plugin for gain calibration. Using the same data and plugin now it gives different results: less hits in the outer layers, compared to earlier analysis in spring/summer of 2017. Simon will look into that.
- Lubomir showed results from straight track runs. In total we had ~120M events, among them ~6M straight tracks with high quality. Alex B. will contact Mike Staib to see if that's enough for his Millipede alignment. We may ask for additional time at the end of the run if needed. The projection of the tracks on FCAL (plot at page 1) shows the exclusion of the inner FCAL blocks (4 on each side, making box 40x40 cm) visible on the plot. What is not understood are the four hot FCAL blocks inside that box, that were supposed to be excluded from the trigger. The wire residuals (page 3) will be compared with previous straight track runs in 2016.
- Alex B. is in contact with Mike and wants to further improve the chamber alignment using Mike's code.
- HV scan runs were analyzed by Lubomir using his straight track code limiting the analysis to small (<6deg) angles where the tracks are almost straight. The hit efficiency vs HV shows that with the nominal HV of 2225V we are well in the plateau (page 4). For the high intensity studies we should stay the DAQ limit and lower the HV till the current in the chamber is acceptable. As shown on the last plot, the loss of efficiency will not be significant.
- Alex A. also analyzed the scan runs with his plugin and showed that the loss of the efficiency for lower HV happens first at bigger drift distances (see plot attached).
- Lubomir discussed the investigation of the HV sector in package 3 cell 4, that doesn't hold HV. By measuring the currents of both positive and negative HV after applying separately positive and negative voltage up to 50V we see that there's an ohmic resistance of 0.5MOhm on the positive and 1MOhm on the negative HV. Each wire has 1MOhm at the HV side, thus the 0.5Mohm can be explained by a broken wire that touches another sense wire and goes to ground on the strips. However the negative HV sector has a common 10KOhm and it is difficult to explain the 1MOhm seen on the negative HV. More studies will be done after the run.
- Next meeting on Feb. 15, will discuss the collaboration meeting talks.